Write an equation in standard form for the line that passes through (-5,-8) and is perpendicular to 3x+2y=-7
Answer by nerdybill(7384)

Write an equation in standard form for the line that passes through (-5,-8) and is perpendicular to 3x+2y=-7
.
slope of:
3x+2y=-7
2y=-3x-7
y=(-3/2)x-7/2
slope is -3/2
a line perpendicular must have a slope that is the negative reciprocal:
(-3/2)m = -1
m = 2/3 (our new slope)
.
Plug 2/3 and (-5,-8) into "point-slope" form:
y - y1 = m(x -x1)
y - (-8) = (2/3)(x -(-5))
y + 8 = (2/3)(x+5)
y + 8 = (2/3)x+10/3
y = (2/3)x+10/3 - 8
y = (2/3)x+10/3 - 24/3
y = (2/3)x - 14/3
that was the "slope-intercept" form, but they wanted in "standard form":
-(2/3)x + y = -14/3 (this is what they're looking for)